a) What output mix should a profit-maximazing firm produce when its total profit function is

πœ‹ = 40π‘₯ βˆ’ 2π‘₯ ΰ¬Ά βˆ’ π‘₯𝑦 βˆ’ 3𝑦 ΰ¬Ά + 100𝑦 and its maximum output capacity is π‘₯ + 𝑦 = 15 ?

b) Estimate the effect on profits if output capacity is expanded by 1 unit.Β 


1
Expert's answer
2021-02-10T21:42:35-0500

Solution

a) What output mix should a profit-maximazing firm produce

πœ‹ = 40π‘₯ βˆ’ 2π‘₯2 βˆ’ π‘₯𝑦 βˆ’ 3𝑦2 + 100𝑦

s.t π‘₯ + 𝑦 = 15


Lang. max = "40𝑥 − 2𝑥^2 − 𝑥𝑦 − 3𝑦^2 + 100𝑦+\lambda(15-x-y)"

"\frac{\delta L}{\delta x}=0 \implies 40-4x-y=\lambda---------(1)\\
\frac{\delta L}{\delta x}=0 \implies -x-6y+100=\lambda--------(2)\\
\frac{\delta L}{\delta x}= 0 \implies x+y=15------------(3)"


Solving (1) and (2) yields

"40-4x-y=-x-6y+100\\
5y-3x=60-----------------(4)"

Solving (3) and (4) yields

"5(15-x)-3x=60 \implies 75-5x-3x=60\\
\therefore -8x=60-75\\
-8x=-15\\
x=1.875\\
y=15-1.875\\
y=13.125"


Hence total profit is;

"\pi=40(1.875)+2(1.875)^2-(1.875)(13.125)-3(13.125)^2+100(13.125)\\
\pi=853.125"


b) Estimate the effect on profits if output capacity is expanded by 1 unit.

"x+y=16------------------(5)"

Solving (4) and (5) yields

"5(16-x)-3x=60 \implies 80-5x-3x=60\\
\therefore -8x=60-80\\
-8x=-20\\
x=2.5\\
y=15-2.5\\
y=12.5"

Hence the total profit in this case is;

"\pi=40(2.5)+2(2.5)^2-(2.5)(12.5)-3(12.5)^2+100(12.5)\\
\pi=862.5"